OpenStudy (anonymous):

How do I graph y=3x+3? This is for a homework assignment but I don't quite understand how to do it.

OpenStudy (anonymous):

Please???

OpenStudy (phi):

y=3x+3 is the equation of a line one way is to find two points on the line, and then draw a line through those two points.

OpenStudy (ashes.boom):

Start with the y intercept with in this case is +3

OpenStudy (ashes.boom):

Then use the slope to find the next point so up 3 over 1

OpenStudy (phi):

to find a point, try x=0 can you find y when x is 0 ?

OpenStudy (phi):

to find y, when x is 0, erase the x and put in zero y = 3*0 + 3 what is 3 time 0 ? and then add 3 ?

OpenStudy (anonymous):

I don't get it sorry im not that great at math. can you explain it a different way?

OpenStudy (anonymous):

would y be 9

OpenStudy (phi):

what is 3 times 0 ?

OpenStudy (anonymous):

0

OpenStudy (phi):

so we start with y =3x+3 put in 0 for x: y = 3*0 + 3 do the operations: remember order of operations, first do the multiply 3*0=0 as you know, so we have y = 0+3 what is y ?

OpenStudy (anonymous):

I think it is 3.

OpenStudy (phi):

yes y=0+3 and 0+3 is 3, so y=3 all of that work means (0,3) is a point on the line. now we find another point. try x=1 in y= 3x+3 can you try ?

OpenStudy (anonymous):

y = 6.

OpenStudy (anonymous):

I think I got it now.

OpenStudy (phi):

so (1,6) is another point. do you know how to plot the points (0,3) and (1,6) ?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

OK thanks

OpenStudy (anonymous):

Thanks Phi and Ashes.Boom for helping.